### v1

# GDI-Bench: A Benchmark for General Document Intelligence with Vision and Reasoning Decoupling

Abstract
The rapid advancement of multimodal large language models (MLLMs) has profoundly impacted the document domain, creating a wide array of application scenarios. This progress highlights the need for a comprehensive benchmark to evaluate these models’ capabilities across various document-specific tasks. However, existing benchmarks often fail to locate specific model weaknesses or guide systematic improvements. To bridge this gap, we introduce a General Document Intelligence Benchmark (GDI-Bench), featuring 1.9k images across 9 key scenarios and 19 document-specific tasks. By decoupling visual complexity and reasoning complexity, the GDI-Bench structures graded tasks that allow performance assessment by difficulty, aiding in model weakness identification and optimization guidance. We evaluate the GDI-Bench on various open-source and closed-source models, conducting decoupled analyses in the visual and reasoning domains. For instance, the GPT-4o model excels in reasoning tasks but exhibits limitations in visual capabilities. To address the diverse tasks and domains in the GDI-Bench, we propose a GDI Model that mitigates the issue of catastrophic forgetting during the supervised fine-tuning (SFT) process through a intelligence-preserving training strategy. Our model achieves state-of-the-art performance on previous benchmarks and the GDI-Bench. Both our benchmark and model will be open source.
[…]
1 Introduction
Rapid progress of large language models (LLMs) [1, 2] has placed multimodal large language models (MLLMs) [3, 4, 5] as a foundation of artificial intelligence, advancing document intelligence to a general stage. Cross-domain and multi-scale document understanding and extraction challenges are increasingly critical in real-world applications. With the rise of MLLMs, a series of more complex benchmarks have emerged to provide comprehensive frameworks for document understanding tasks [6, 7, 8, 9, 10, 11, 12]. However, given that document understanding involves multiple modalities, errors in MLLM outputs may arise from inaccurate visual recognition, limited language organization, or both. Consequently, a decoupled evaluation of MLLMs’ document processing abilities is essential.
To address these challenges, we propose the General Document Intelligence Benchmark (GDI-Bench), which aims to perform a decoupled evaluation of model performance in document tasks, thereby contributing to the identification of the model’s weaknesses. The GDI-Bench introduces three key improvements over existing benchmarks: (1) developing a cross-domain, multi-task benchmarks to ensure task diversity and fine-grained difficulty levels; (2) introducing complexity decoupling, dividing multimodal document understanding into visual complexity and reasoning complexity, and for the first time, establishing a difficulty grading mechanism; (3) supporting the evaluation of MLLMs, OCR+LLM-level systems, and document parsing tools, offering comprehensive guidance for practical application solutions.
We evaluate 2 open-source and 4 closed-source large models using the GDI-Bench and find their performance to be suboptimal. To enhance the model’s performance on GDI-Bench, we further conduct supervised fine-tuning (SFT) on the InternVL3-8B model to explore whether a data-driven approach could help the model acquire extensive domain-specific document knowledge. However, we observe that supervised fine-tuning tends to cause the issue of catastrophic forgetting. To address this problem, we propose the Layer-wise Adaptive Freezing Tuning (LW-AFT) method, which alleviates the impact of catastrophic forgetting and enhances the model’s cross-domain and cross-task capabilities. Specifically, during the SFT process, LW-AFT freezes most of the parameters, with only a small subset of domain-sensitive parameters participating in gradient updates.
Our contributions are summarized as follows.
  • We propose the GDI-Bench, which encompasses a broad range of tasks in the document domain. By decoupling complexity and grading difficulty, it helps the model identify its weaknesses and guides subsequent optimization.
  • We propose Layer-wise Adaptive Freeze-Tuning, a training method that effectively alleviates catastrophic forgetting in document task SFT by parameter-freezing, improving performance on specific tasks while maintaining generalization capabilities.
  • We propose a GDI model that achieves state-of-the-art (SOTA) performance on multiple document domain benchmarks as well as the GDI-Bench, demonstrating high generalization capabilities suitable for real-world applications.
2 Related Works
2.1 Document Benchmark
Early benchmarks in document understanding predominantly targeted single-domain data and specific tasks. For instance, DocVQA[13] focused on industrial document QA, VisualMRC[14] addressed web-based document comprehension, and ChartQA[15] specialized in chart-based question answering. Despite their contributions, these benchmarks lacked difficulty stratification and cross-domain generalization, largely due to the limited generalization capabilities of models available at the time. With the rise of multimodal large models, a series of more complex benchmarks emerged [6, 7, 8, 9, 10, 11, 12], aiming to provide a more comprehensive framework for document understanding tasks. OCRBenchV2[16] provides multiple OCR-related subtasks, expanding the scope of OCR evaluation. Fox[17] introduced region-of-interest (ROI)-based document parsing methods, enhancing the detailed analysis of document structures, while OmniDocBench[18] extended the evaluation to cross-modal tasks, covering a broader range of document parsing challenges. While these benchmarks have broadened the scope of evaluation, they still lack explicit difficulty grading, which limits their ability to assess model performance across tasks of varying complexity.
2.2 Document Understanding Model
Optical Character Recognition (OCR) has long been a fundamental task in computer vision. Existing OCR models can be broadly categorized into component-based and end-to-end approaches. Component-based methods [19, 20, 21] adopt a modular pipeline that assembles multiple expert-designed components such as layout analysis [22], text detection [23, 24, 25, 26], region extraction, and content recognition [27, 28, 29]. In contrast, end-to-end OCR models [30], especially those driven by Large Vision-Language Models (LVLMs)[31, 3, 32, 33, 34, 35, 36], aim to unify perception and reasoning within a single architecture. Most LVLM-based OCR systems utilize CLIP [37] as the vision backbone, while coupling it with a language model to jointly process visual and textual information in a unified framework. Recent works [34, 35, 38] adopt sliding-window strategies that partition the image into patches to cope with long and high-resolution inputs like PDFs.
2.3 Continual Learning of LLMs
Continual learning in large language models (LLMs) faces the core challenge of catastrophic forgetting. Existing research mainly adopts two strategies: data replay  [39, 40] and parameter freezing [41, 42, 43]. Data replay strategies mitigate forgetting by revisiting samples from previous tasks during the training of new tasks. A typical approach is LAMOL[39], which generates pseudo-samples using the language models to avoid storage-based replay. Other approaches, such as experience replay and interleaved task training, have also been used to incorporate old data into the training loop. The parameter freezing method, on the other hand, protects existing knowledge by restricting parameter updates, such as through fine-tuning lightweight modules like LoRA[44], adapters [45], or prompt tokens [46]. Classical regularization-based methods like Elastic Weight Consolidation (EWC)  [47] protect crucial weight parameters through importance constraints. More recent methods, such as Task Vector [48] and Gradient Projection [49], operate on the gradient or representation level to minimize interference between tasks and preserve generalization. In contrast to these methods, which generally rely on a limited number of pre-training datasets, leading to incomplete domain coverage and significant computational costs, our method presents a more comprehensive and efficient alternative.
3 Benchmark
Figure 1: The overview of GDI-Bench.
3.1 Complexity Decoupling
As illustrated in Figure 1, a difficulty-decoupled evaluation protocol named GDI-Benchmark is introduced to comprehensively assess MLLMs’ capabilities in visual information comprehension and reasoning. Distinct from VisualSimpleQA[50]’s structural decomposition of fact-seeking question answering tasks into visual recognition and knowledge dimensions through VLLM model architecture analysis, the proposed framework characterizes task complexity from fundamental difficulty perspectives by decoupling it into visual complexity and reasoning complexity components. Notably, knowledge popularity is explicitly subsumed within the reasoning complexity dimension in this formulation.
3.1.1 Visual Complexity
Figure 2: Distribution of visual complexity scores.
The visual complexity dimension is operationalized through a hierarchical categorization of document images into three levels: V0 (plain text), V1 (formal representations), and V2 (explanatory representations). V0 exclusively contains unstructured textual elements such as headings and paragraphs. Multimodal tasks on V0 documents typically achieve satisfactory performance via OCR-LLM pipeline architectures. As demonstrated in Figure 2, systematic analysis of the Omnidocbench benchmark reveals statistically significant performance gaps. This benchmark covers nine document categories. Current pipeline tools and VLLMs show notably worse performance on textbook, exam paper, notes, and newspaper types compared to other forms. These performance drops directly correlate with complex features like multi-column layouts and graphical elements. Based on these observations, we designate these four challenging categories as V2 and classify the remaining five as V1. This creates a data-driven taxonomy for visual complexity characterization.
3.1.2 Reasoning Complexity
The reasoning complexity characterization is formulated through a behavior-driven taxonomy. Three distinct levels are defined to progressively evaluate document understanding capabilities. It is specifically categorized into R0: Full Page Structured Extract, R1: Information Extract, and R2: Reasoning. Among these, the R0 task in V0 is more similar to OCR, and requires the ability to understand layout information such as tables, formulas, and complex page structures in V1 and V2 type images. For R1 tasks, the model needs not only to recognize visual content but also to understand the task itself in order to accurately extract relevant information, this often involves interpreting lines, tables, labels, and other visual elements present in V1 and V2 images. R2 tasks go a step further by requiring deeper reasoning capabilities, including the comprehension of logical inference and the ability to synthesize information across different modalities or layouts.
3.2 Annotation Process
Figure 3: Annotation Process of GDI-Bench.
3.2.1 Data Source
For the data acquisition phase of the General Document Intelligence (GDI) Benchmark, documents were primarily sourced from Omnidocbench[18], a comprehensive dataset containing document images from 9 domains. The dataset was further supplemented with an in-house collection of various document types, including exam papers, reports, newspapers and so on. This multi-source integration resulted in a well-rounded and representative dataset that captures the multifaceted nature of real-world document understanding tasks.
3.2.2 Data Construction
As shown in Figure 3, the data construction process begins with cropping single-layout sub-images from Omnidocbench[18] and inhouse documents to form the V0 raw image set. Based on end-to-end edit distance scores from state-of-the-art models and pipeline tools on Omnidocbench, domains with scores above 0.142 are categorized as V2, indicating high visual complexity. The remaining samples are assigned to V1.
For task construction, R0 tasks are generated using the original annotations or by synthesizing Markdown representations through MinerU[19]. To create R1 and R2 tasks, both the Markdown and corresponding images are input into GPT-4o to generate extractive and reasoning-based question-answer pairs. In addition, rule-based tasks are designed manually. These include extracting questions with fixed numbering, such as identifying problem numbers in exam papers.
To ensure high data quality, all synthetic cases are first evaluated by models. Cases that can be answered correctly by DeepSeek without requiring document input or that are flagged as low quality are filtered out. Finally, a team of PhD-level annotators reviews and verifies the remaining instances to ensure accuracy and correctness in the final benchmark dataset.
The final GDI-Benchmark contains a total of 2,989 test cases, distributed across different visual complexity levels (V0, V1, V2) and task types (R0, R1, R2), as summarized in Table 1.
Table 1: Complexity Distribution in the GDI Benchmark
v0 v1 v2 Total
r0 379 528 453 1 360
r1 223 407 276 906
r2 133 311 279 723
Total 735 1 246 1 008 2 989
3.3 Evaluation Metrics
Metrics are defined per task type. R0 uses Average Edit Distance (AED) for character-level recognition accuracy. R1 adopts Average Normalized Levenshtein Similarity (ANLS) for entity and field extraction evaluation. For R2, Accuracy is the primary metric, as most tasks require single-choice answers. A small number of structured reasoning tasks in R2 require models to generate structured outputs, for which AED is used to assess similarity to reference answers.
3.4 Comparison with Existing Document Intelligence Benchmarks
Table 2: Comparison of GDI-Bench with Other Document Understanding Benchmarks
Benchmark Scenario Task Image Difficulty Grading
Seed-bench-2-plus  8 1 0.6k
MMTab-eval 1 9 23k
MMC 1 9 1.7k
OCRBenchV2 31 23 9.5k
Fox 2 9 0.7k
GDI-Bench 9 19 1.9k
As shown in Table 2, most existing document intelligence benchmarks are limited in either scenario coverage, task diversity, or data scale. In contrast, the GDI-Bench offers a balanced combination of 9 diverse document scenarios and 19 representative tasks, based on a curated set of 1.9k images. Notably, it introduces difficulty grading , a feature absent in all other compared benchmarks.
Table 3: Composition of the training set.
Domain Task Training data
Newspaper
Full text extraction of sliced newspaper pages.
Extract header information of newspaper pages.……
29k
Scientific Paper
Extract the author information of the paper.
Extract full-text page content and organize it in markdown format.
Answer reasoning questions for a paragraph on a page scientific paper page (/ multiple choice).……
108k
Infograph
Extract the maintitle.
Answer reasoning questions for a infographic page (/ multiple choice).……
32k
Exam paper
Full text extraction of exampaper pages.
Extract the content of the exam paper based on the type of question and the question number.……
12k
Financial report
Extract the table on a page and convert to LaTeX format.
Extract the content of the exam paper based on the type of question and the question number.
3k
Handwritten Content Recognize handwritten content in English and Chinese. 12k
4 Methodology
(a) The mean values of parameter changes per layer before and after full-parameter SFT.
(b) Distribution of parameter differences before and after full-parameter SFT.
GDI-Bench presents unique challenges for model designs, motivating the application of supervised fine-tuning to improve task-specific performance. To tackle this, we develope a multi-source training set that includes tasks structurally aligned with the benchmark, but drawn from diverse data domains, as shown in Table 3. We perform SFT on the InternVL3-8B model using this training set. However, subsequent evaluation of the model reveals significant catastrophic forgetting, wherein the model loses essential knowledge acquired during pre-training.
To investigate the phenomenon, we conduct a parameter update analysis of models undergoing full-parameter supervised fine-tuning. Our experimental results reveal significant sparsity in learned parameter adaptations: over 95% of parameters exhibit minimal changes, while a critical sparse subset (5%) undergoes substantial modification (>0.005), as depicted in Fig. 4(b). This finding aligns with and extends the Lottery Ticket Hypothesis [51], leading to our formal hypothesis: For any pre-trained MLLM with parameters $\Theta=\{\theta_{i}\}_{i=1}^{n}$, there exists a sparse task-salient sub-network $\hat{\Theta}=\{\hat{\theta}_{j}\}_{j=1}^{m}$ ($m<n$) and a domain-specific transformation $\psi$, such that $\psi(f_{\hat{\Theta}}(x))$ achieves performance comparable to the full model $f_{\Theta}(x)$ on the target domain $\mathcal{D}$.
$\forall x\in\mathcal{D},\ \mathcal{P}(f_{\Theta}(x))\approx\mathcal{P}(\psi(f_{\hat{\Theta}}(x)))$ (1)
where $\mathcal{P}(\cdot)$ is the metric for evaluating the performance of the model, and the subnetwork $\hat{\Theta}$ consists of parameters exhibiting high update sensitivity during SFT. Since only a small subset of parameters is relevant to task-specific adaptation, we propose performing gradient updates solely on these parameters during SFT, while freezing the majority of the remaining parameters to preserve the model’s generalization and essential knowledge as much as possible.
We formalize this approach as Layer-wise Adaptive Freeze-Tuning (LW-AFT), which identifies and updates only critical parameters through layer-wise sensitivity analysis. By maintaining stable base network dynamics and avoiding large-scale parameter shifts, LW-AFT mitigates catastrophic forgetting while enabling efficient adaptation to downstream tasks. The overview of our method is shown in Fig. 5. For a specific new domain $\mathcal{D}_{t}$, the domain-specific expert model (denoted as $\mathcal{E}_{t}$) is fine-tuned based on the pre-trained parameters $\Theta$, using only a fraction $\frac{1}{\alpha}$ of the full dataset. For each domain-specific expert model, the element-wise absolute change in parameters is computed as:
$\Delta\theta^{(t)}_{i}=\left|\theta^{\mathcal{E}_{t}}_{i}-\theta_{i}\right|.$ (2)
We perform a comprehensive layer-wise analysis of parameter updates $\Delta\theta^{(t)}_{i}$ before and after SFT. Fig. 4(a) illustrates the average magnitude of these updates per layer, revealing distinct trends between the vision and language components of the model. Specifically, vision layers exhibit more parameter modifications compared to their language layers. Intuitively, we assume that the layers undergoing substantial parameter updates are critical for domain adaptation, as their dynamics closely align with task-specific knowledge transfer. Conversely, layers with minimal updates appear less specialized to domain-specific features, and preserving their stability supports cross-domain generalization.
Figure 5: Overview of the Layer-wise Adaptive Freeze-Tuning method.
Therefore, we implement a layer-adaptive unfrozen parameter allocation strategy. Partitioning the model into $L$ architectural layers with parameter counts $W=\{w_{1},...,w_{L}\}\subseteq\mathbb{R}$, the distribution of trainable parameters across layers follows a proportional allocation mechanism: Given a global budget of $H$ unfrozen parameters, each layer $\ell$ receives a number of unfrozen parameters $h_{\ell}$ determined by
$h_{\ell}=\frac{m_{\ell}\cdot w_{\ell}}{\sum_{i=1}^{L}m_{i}\cdot w_{i}}\cdot H$ (3)
where $m_{\ell}=\text{Avg}(\{\Delta\theta_{j}^{(t)}\}_{\ell})$ is the average absolute change of the $\ell$-th layer.
After determining the number of unfreezing parameters for each layer $\ell$, we first sort the absolute changes $\{\Delta\theta_{j}^{(t)}\}_{\ell}$ in descending order, and then select only the top $h_{\ell}$ parameters for further updates. This selection is performed via the function:
[…]
The binary masks obtained for the expert $\mathcal{E}_{t}$ is denoted as $\mathcal{M}_{\mathcal{E}_{t}}=\left\{\phi_{h_{\ell}}\bigl{(}\{\Delta\theta_{j}^{(t)}\}_{\ell}\bigr{)}\right\}_{\ell=1}^{L}$. In the subsequent fine-tuning on a target domain, the gradient update is modified so that only the parameters with $\mathcal{M}_{\mathcal{E}_{t}}$ are updated:
$\nabla_{\theta}\mathcal{L}(\theta)\leftarrow\mathcal{M}_{\mathcal{E}_{t}}\odot\nabla_{\theta}\mathcal{L}(\theta),$ (5)
where $\odot$ denotes the Hadamard product. This approach not only reduces the number of parameters to be trained but also preserves full-domain generalization without resorting to experience replay or other sophisticated fine-tuning techniques.
5 Experiments
The experiments in this section consist of two parts. In Section 5.1, we discuss the impact of our proposed Layer-wise Adaptive Freeze-Tuning (LW-AFT) method on cross-domain and cross-task generalization. In Section 5.2, we perform a comprehensive benchmark evaluation on the GDI-Bench.
Table 4: The performance of different training methods on cross-domain and cross-task scenarios.
Model T1 T2 T3 T4
date editor email phone
InternVL 3-8B 0.216 0.383 0.840 0.101 0.884 0.398 0.398
Full-Parameter Fine-Tuning 0.040 0.913 0.600 0.864 0.992 0.783 0.989
LoRA Fine-Tuning 0.102 0.473 0.300 0.093 0.606 0.171 0.172
LW-AFT (Ours) 0.096 0.365 0.200 0.010 0.316 0.058 0.153
5.1 Layer-wise Adaptive Freeze-Tuning.
5.1.1 Cross-domain and cross-task evaluation.
To validate the generalization capability of the LW-AFT method in cross-domain and cross-task scenarios, we designed the following experiments.
Cross-domain experiments with the same task.
  • T1: Train the paragraph starting position OCR task in the academic paper domain and test the same task in the Infograph domain.
  • T2: Train the information formatting and organization (JSON) task in the exam and academic paper domains, and test it in the Infograph domain.
  • T3: Train the reasoning and question-answering task in the academic paper domains, and test it in the financial report domain.
Cross-task experiments with the same domain.
  • T4: In the newspaper domain, only train the header information extraction and its formatting organization (JSON) task, and test the extraction task for header detail information (including extraction of newspaper date, editor, email, and phone number).
We compare the performance of full-parameter fine-tuning, LoRA fine-tuning, and our method under both settings, as shown in Table 4, which displays the edit distance for each task. As can be seen, our method demonstrates strong cross-domain and cross-task capabilities, significantly outperforming the LoRA fine-tuning.
5.1.2 Performance on general datasets.
We evaluate the performance of full-parameter fine-tuning, LoRA fine-tuning, and our method on a range of general datasets, as shown in Table 5. It can be seen that our method preserved the model’s general capabilities and avoided catastrophic forgetting after SFT.
Table 5: The performance of different training methods on general datasets.
[…]
val test test no mask test - val - en cn
InternVL 3 8B 92.0 86.6 85.2 92.6 80.2 880 75.6 69.7 85.5 85.6
Full-Parameter Fine-Tuning 39.4 33.3 43.1 49.4 22.7 228 25.9 44.8 15.6 14.6
LoRA Fine-Tuning 91.3 85.8 85.4 93.1 81.6 862 74.4 69.7 85.2 84.8
LW-AFT (Ours) 91.6 85.8 85.6 93.2 82.1 871 75.8 70.6 85.2 84.8
Table 6: The performance of different training methods on general datasets.
Model R0V0 R0V1 R0V2 R1V0 R1V1 R1V2 R2V0 R2V1 R2V2 Overall
Claude3-5V-Sonnet 0.10 0.74 0.78 0.60 0.52 0.91 0.18 0.52 0.62 0.58
Gemini-2.0-Flash 0.05 0.42 0.52 0.62 0.25 0.55 0.05 0.31 0.69 0.39
GPT4o-2024-11-20 0.04 0.40 0.57 0.59 0.40 0.58 0.12 0.48 0.68 0.43
Grok-2-Vision 0.33 0.80 0.84 0.80 0.71 0.70 0.23 0.54 0.62 0.66
Qwen2.5-VL-72B-Instruct 0.03 0.35 0.44 0.41 0.24 0.33 0.11 0.36 0.50 0.31
GDI-Model 0.03 0.35 0.44 0.14 0.21 0.24 0.08 0.23 0.34 0.25
5.2 Benchmark Evaluation Results.
To assess the effectiveness of our GDI model on generalized OCR tasks, we compare it against several state-of-the-art vision-language models—namely Qwen2.5-VL-72B, Gemini-2.0-Flash, GPT-4o-2024-11-20, Claude-3-5-Sonnet, and Grok-2-Vision—on GDI-Bench’s multi-level challenge suite (v0-r0 through v2-r2). All evaluations were conducted with the same preprocessing pipeline.
On v0-r0, our 8B model matches or slightly exceeds performance of Qwen2.5-VL-72B. On the held-out generalization set, GDI-Model retains a slight advantage over the much larger models (70 B+), demonstrating superior resource-efficiency and robustness. These results confirm that, despite its smaller parameter count, our model achieves performance on par with or exceeding that of much larger open- and closed-source competitors.
6 Conclusion
We introduce the GDI-Bench, a document-domain benchmark with extensive domain coverage that pioneers a systematic difficulty grading system. Utilizing a complexity decoupling mechanism, the GDI-Bench breaks down multimodal document tasks into two orthogonal dimensions: visual complexity and reasoning complexity. It establishes a quantifiable, hierarchical standard that aids in identifying model issues and guiding optimization efforts. In addition, we propose the Layer-wise Adaptive Freeze-Tuning method and the corresponding GDI model. LW-AFT mitigates the problem of catastrophic forgetting during SFT, retains 99% of the model’s general capabilities during the SFT process, and enhances its performance on cross-domain and cross-task scenarios. The GDI model has demonstrated outstanding performance across multiple document-domain benchmarks, including the GDI-Bench, as well as general datasets. We hope that the GDI-Bench can help the advancement of future models and perhaps inspire new theoretical insights.